# From Adversarial Poetry to Adversarial Tales: An Interpretability Research Agenda

### v1

@@ 2 Related Work
Empirical evidence has shown that aligned language models remain susceptible to adversarial manipulation, despite extensive efforts to increase their safety measures. Jailbreaking techniques exploit this gap through deliberate input crafting designed to circumvent safety, ethical or legal constraints. Existing taxonomies classify jailbreaks by strategy and their underlying mechanism rao2024trickingllmsdisobedienceformalizing; schulhoff2023hackaprompt.
Common strategies include structural perturbations of the malicious request, the construction of fictional or virtual scenarios kang2023exploitingprogrammaticbehaviorllms, exploitation of the tendency of the model to maintain helpfulness and coherence rao2024trickingllmsdisobedienceformalizing; perez2022ignorepreviouspromptattack (sometimes prioritizing them over safety boundaries), persuasive appeals with urgency cues zeng2024johnnypersuadellmsjailbreak and feedback-driven exploitation of distributional weaknesses zou2023universal; in practice, effective attacks frequently combine multiple strategies within a single prompt. Contexts perceived as low-risk, such as stories, games or in general simulated scenarios, increase the likelihood that models bypass safety constraints, potentially because policies are treated as inapplicable within virtual frames.
Other works have, for instance, explored nested virtualization, where the harmful objective is recursively wrapped inside layered fictional scenes that the model is instructed to continue (li2024deepinceptionhypnotizelargelanguage) or the automated synthesis of narrative-based attacks at scale (ntais2025jailbreakmimicryautomateddiscovery). Furthermore, Many of the principles of fictional contexts are explored by successful jailbreak prompts such as the “Do Anything Now” (DAN) family shen2024donowcharacterizingevaluating of attacks.
Recently, hierarchical, multi-step reasoning has been identified as a critical failure mode. Structured analytical tasks can increase ASR by shifting model objectives from refusal toward task completion wang2025safetylargereasoningmodels; zhao2025chainofthoughthijacking. These findings suggest that forcing models into reasoning at multiple levels of abstraction may exploit vulnerabilities beyond surface-level perturbations.
@@ 3.3.1 Vladimir Propp’s Morphology of the Folktale
In his Morphology of the Folktale, Vladimir propp1968morphology introduced a structural approach to narrative analysis that remains influential across folklore studies, literary theory, and computational narratology. Propp observed that Russian folktales, despite their surface diversity, share a common deep structure: a fixed sequence of narrative functions, defined as acts of characters from the standpoint of their significance for the course of action.
@@ 3.3.2 Attack Construction
To illustrate the attack structure without enabling direct reproduction, we provide the following sanitized proxy. The actual harmful content has been replaced with an innocuous procedure (baking a cake), but the structural template is preserved.
@@ 3.4 Experimental Setup
Drawing from bisconti2025adversarial, we measured the effectiveness of the attacks using ASR metric, defined as:
@@ 3.5 Results
Comparing these findings with Adversarial Poetry (bisconti2025adversarial) reveals both continuity and escalation. Poetry achieved 62% overall ASR across 25 models; Tales reaches 71.3% across 26 models. The provider ranking shifts notably: Anthropic, which showed strong resistance to Poetry (average ASR below 35%), now exhibits 47.5% ASR under Tales. OpenAI models, previously the most robust family against Poetry with ASR as low as 0-10% for smaller models, now range from 35% to 57%. Google models remain highly vulnerable under both attacks (86.7% for Tales vs. 65-100% for Poetry). The consistency of vulnerability across both attack types, despite their distinct mechanisms, supports the hypothesis that these are instances of a broader class of structurally-grounded jailbreaks rather than isolated techniques.
Unlike Adversarial Poetry (bisconti2025adversarial), where larger models generally exhibited higher ASR, no substantial size effect emerges for Adversarial Tales. This suggests that the functional analysis framing exploits a more fundamental property of how models process structured interpretive tasks, rather than a capability that scales with model size.
@@ 4 Discussion: Adversarial Poetry and Tales
According to  wei2023jailbrokendoesllmsafety, jailbreaks are built on two main alignment weaknesses of the models, namely Competing Objectives where the model prioritizes goals that conflict with safety rules and Mismatched Generalization, where refusal behavior fails to generalize to out-of-distribution inputs that differ in surface from the ones encountered in safety-training.
@@
The model is still explicitly asked to carry out a harmful task, but the request is encoded in verse and metaphor, inducing a stylistic distribution shift that moves the input outside the safety-training distribution. Poetic prompts occupy regions of the input space abundantly represented during pretraining but underrepresented in alignment data, leading models to prioritize helpfulness and literary interpretation over safety policies. By contrast, Adversarial Tales expose a more complex failure mode in which stylistic shift, fictional framing, hierarchical reasoning and juxtapositions of objectives interact. Following wei2023jailbrokendoesllmsafety, mismatched generalization arises from the culturally coded narrative, while competing objectives emerge from the well-framed and seemingly benign request, so that the model privileges legitimate interpretation over safety, overriding refusal. While we frame Adversarial Poetry as primarily exploiting mismatched generalization and Adversarial Tales as engaging competing objectives, in practice both mechanisms likely operate simultaneously in both attack classes. The cleaner distinction may be that Adversarial Tales layer fictional framing and hierarchical task decomposition onto the stylistic distribution shift already present in poetic reformulation, creating a composite vulnerability surface rather than an entirely distinct failure mode.
@@ 5 Towards New Defenses: A Research Agenda in Mechanistic Interpretability
Broadly speaking, mechanistic interpretability research can be divided into two methodological approaches, as discussed by sharkey2025open: a reverse engineering approach, which aims to identify the functional roles of network components, and a concept-based approach, which aims to identify the network components responsible for given roles. Because working with concepts in the latent space of large LMs is particularly demanding in a resource-constrained setting, we initially concentrate on the reverse engineering approach, which proceeds by decomposing the model into its constituent parts.
A naive way to perform such a decomposition is to analyze neural networks at the level of architectural components, such as individual neurons, activation heads, or layers. The core critique is that individual attention heads may respond to multiple distinct features (polysemanticity) janiak2023polysemantic, and that attention patterns themselves can be misleading as explanations of model behavior jain2019attention. This limitation may become even more pronounced in modern large-scale models, where increased capacity is associated with more complex and often more polysemantic internal features.
Yet in many practical settings, such as detecting and mitigating contextual hallucinations, analyzing the behavior of individual attention heads or of attention patterns as a whole can be instrumental. For example, wu2024retrieval identify a specialized class of attention heads, retrieval heads, that are strongly implicated in hallucinations in long-context settings and in performance on needle-in-a-haystack tasks. Similarly, chuang2024lookback argue that hallucinations are related to the extent to which an LLM attends to information in the provided context versus its own past generations, and on this basis develop an attention-based hallucination detection model, illustrating how analyzing attention can inform the design of new defenses against undesirable LLM behaviors.
@@
Given distinct attentions patterns under narrative attacks, the cross-model generality of this mechanism could be understood in light of kaushik2025universal. They provide evidence that neural networks systematically converge to shared spectral subspaces, regardless of initialization, task, or domain, thereby motivating what they call the universal weight subspace hypothesis. If this hypothesis is correct, the resulting lack of diversity may represent a fundamental bottleneck, implying that models inherit shared biases, capabilities, and failure modes, including vulnerability to Adversarial Poetry and Adversarial Tales attacks. This perspective can help explain why narrative jailbreaks generalize across models and why the associated vulnerability is systemic, and it can also guide the development of mechanistic-interpretability-driven anti-jailbreaking methods that are transferable across architectures.

### v2

@@ 2 Related Work
Empirical evidence has shown that aligned language models remain susceptible to adversarial manipulation, despite extensive efforts to increase their safety measures. Jailbreaking techniques exploit this gap through deliberate input crafting designed to circumvent safety, ethical or legal constraints. Existing taxonomies classify jailbreaks by strategy and their underlying mechanism Rao et al. (2024); Schulhoff et al. (2023).
Common strategies include structural perturbations of the malicious request, the construction of fictional or virtual scenarios Kang et al. (2023), exploitation of the tendency of the model to maintain helpfulness and coherence Rao et al. (2024); Perez and Ribeiro (2022) (sometimes prioritizing them over safety boundaries), persuasive appeals with urgency cues Zeng et al. (2024) and feedback-driven exploitation of distributional weaknesses Zou et al. (2023); in practice, effective attacks frequently combine multiple strategies within a single prompt. Contexts perceived as low-risk, such as stories, games or in general simulated scenarios, increase the likelihood that models bypass safety constraints, potentially because policies are treated as inapplicable within virtual frames.
Other works have, for instance, explored nested virtualization, where the harmful objective is recursively wrapped inside layered fictional scenes that the model is instructed to continue (Li et al., 2024) or the automated synthesis of narrative-based attacks at scale (Ntais, 2025). Furthermore, Many of the principles of fictional contexts are explored by successful jailbreak prompts such as the “Do Anything Now” (DAN) family Shen et al. (2024) of attacks.
Recently, hierarchical, multi-step reasoning has been identified as a critical failure mode. Structured analytical tasks can increase ASR by shifting model objectives from refusal toward task completion Wang et al. (2025); Zhao et al. (2025). These findings suggest that forcing models into reasoning at multiple levels of abstraction may exploit vulnerabilities beyond surface-level perturbations.
@@ 3.3.1 Vladimir Propp’s Morphology of the Folktale
In his Morphology of the Folktale, Vladimir Propp (1968) introduced a structural approach to narrative analysis that remains influential across folklore studies, literary theory, and computational narratology. Propp observed that Russian folktales, despite their surface diversity, share a common deep structure: a fixed sequence of narrative functions, defined as acts of characters from the standpoint of their significance for the course of action.
@@ 3.3.2 Attack Construction
To illustrate the attack structure without enabling direct reproduction, we provide the following sanitized proxy. The actual harmful content has been censored, preserving the structural template.
@@ 3.4 Experimental Setup
Drawing from Bisconti et al. (2025), we measured the effectiveness of the attacks using ASR metric, defined as:
@@ 3.5 Results
Comparing these findings with Adversarial Poetry (Bisconti et al., 2025) reveals both continuity and escalation. Poetry achieved 62% overall ASR across 25 models; Tales reaches 71.3% across 26 models. The provider ranking shifts notably: Anthropic, which showed strong resistance to Poetry (average ASR below 35%), now exhibits 47.5% ASR under Tales. OpenAI models, previously the most robust family against Poetry with ASR as low as 0-10% for smaller models, now range from 35% to 57%. Google models remain highly vulnerable under both attacks (86.7% for Tales vs. 65-100% for Poetry). The consistency of vulnerability across both attack types, despite their distinct mechanisms, supports the hypothesis that these are instances of a broader class of structurally-grounded jailbreaks rather than isolated techniques.
Unlike Adversarial Poetry (Bisconti et al., 2025), where larger models generally exhibited higher ASR, no substantial size effect emerges for Adversarial Tales. This suggests that the functional analysis framing exploits a more fundamental property of how models process structured interpretive tasks, rather than a capability that scales with model size.
@@ 4 Discussion: Adversarial Poetry and Tales
According to  Wei et al. (2023), jailbreaks are built on two main alignment weaknesses of the models, namely Competing Objectives where the model prioritizes goals that conflict with safety rules and Mismatched Generalization, where refusal behavior fails to generalize to out-of-distribution inputs that differ in surface from the ones encountered in safety-training.
@@
The model is still explicitly asked to carry out a harmful task, but the request is encoded in verse and metaphor, inducing a stylistic distribution shift that moves the input outside the safety-training distribution. Poetic prompts occupy regions of the input space abundantly represented during pretraining but underrepresented in alignment data, leading models to prioritize helpfulness and literary interpretation over safety policies. By contrast, Adversarial Tales expose a more complex failure mode in which stylistic shift, fictional framing, hierarchical reasoning and juxtapositions of objectives interact. Following Wei et al. (2023), mismatched generalization arises from the culturally coded narrative, while competing objectives emerge from the well-framed and seemingly benign request, so that the model privileges legitimate interpretation over safety, overriding refusal. While we frame Adversarial Poetry as primarily exploiting mismatched generalization and Adversarial Tales as engaging competing objectives, in practice both mechanisms likely operate simultaneously in both attack classes. The cleaner distinction may be that Adversarial Tales layer fictional framing and hierarchical task decomposition onto the stylistic distribution shift already present in poetic reformulation, creating a composite vulnerability surface rather than an entirely distinct failure mode.
@@ 5 Towards New Defenses: A Research Agenda in Mechanistic Interpretability
Broadly speaking, mechanistic interpretability research can be divided into two methodological approaches, as discussed by Sharkey et al. (2025): a reverse engineering approach, which aims to identify the functional roles of network components, and a concept-based approach, which aims to identify the network components responsible for given roles. Because working with concepts in the latent space of large LMs is particularly demanding in a resource-constrained setting, we initially concentrate on the reverse engineering approach, which proceeds by decomposing the model into its constituent parts.
A naive way to perform such a decomposition is to analyze neural networks at the level of architectural components, such as individual neurons, activation heads, or layers. The core critique is that individual attention heads may respond to multiple distinct features (polysemanticity) Janiak et al. (2023), and that attention patterns themselves can be misleading as explanations of model behavior Jain and Wallace (2019). This limitation may become even more pronounced in modern large-scale models, where increased capacity is associated with more complex and often more polysemantic internal features.
Yet in many practical settings, such as detecting and mitigating contextual hallucinations, analyzing the behavior of individual attention heads or of attention patterns as a whole can be instrumental. For example, Wu et al. (2024) identify a specialized class of attention heads, retrieval heads, that are strongly implicated in hallucinations in long-context settings and in performance on needle-in-a-haystack tasks. Similarly, Chuang et al. (2024) argue that hallucinations are related to the extent to which an LLM attends to information in the provided context versus its own past generations, and on this basis develop an attention-based hallucination detection model, illustrating how analyzing attention can inform the design of new defenses against undesirable LLM behaviors.
@@
Given distinct attentions patterns under narrative attacks, the cross-model generality of this mechanism could be understood in light of Kaushik et al. (2025). They provide evidence that neural networks systematically converge to shared spectral subspaces, regardless of initialization, task, or domain, thereby motivating what they call the universal weight subspace hypothesis. If this hypothesis is correct, the resulting lack of diversity may represent a fundamental bottleneck, implying that models inherit shared biases, capabilities, and failure modes, including vulnerability to Adversarial Poetry and Adversarial Tales attacks. This perspective can help explain why narrative jailbreaks generalize across models and why the associated vulnerability is systemic, and it can also guide the development of mechanistic-interpretability-driven anti-jailbreaking methods that are transferable across architectures.